The goal of this "de-tuning" run plan is to try to regenerate (briefly!) the large background/lumi detector asymmetries observed in: https://hallcweb.jlab.org/hclog/1112_archive/111211121719.html
The idea is to adjust back the three focussing elements that Poelker/Yan changed on Friday to try to regenerate the large charge asymmetry widths and background asymmetries.
They will back out in three steps; we will take short 5 minute runs at each; then a longer (~ 1 hour run). The back-out steps are:
* Remove the focussing element tweak that affected A4/A3 transmission
* In two steps, remove the focussing element tweaks that affected the A2 transmission.